OpenStudy (anonymous):

6. Solve: 4x + 2 (x - 6) = 30 Step 1: 4x + 2x - 6 = 30 Step 2: 6x - 6 = 30 Step 3: 6x = 30 Step 4: x = 5 Which is the first incorrect step in the solution shown above? A. Step 1 C. Step 3 B. Step 2 D. Step 4

OpenStudy (anonymous):

A step 1

OpenStudy (anonymous):

4x+2x-12 not 6

OpenStudy (anonymous):

If you know how to apply the distributive property you'll know that it's the first one.

OpenStudy (anonymous):

what is the answer? I'm getting different ones?

OpenStudy (anonymous):

are u sure? :)

OpenStudy (anonymous):

So what step? ^.^

OpenStudy (anonymous):

But 6 times 5 is 30. so it can't be step 4

OpenStudy (anonymous):

8. Solve: 11x - 9 = 5x - 12x + 19 Step 1: 11x - 9 = -7x + 28 Step 2: 11x = -7x + 28 Step 3: 4x = 28 Step 4: x = 7 Which is the first incorrect step in the solution shown above? A. Step 1 C. Step 3 B. Step 2 D. Step 4

OpenStudy (anonymous):

@Luis Rivera

OpenStudy (anonymous):

@Luis Rivera